An electric heater which produces 900W of power is used to vaporize water. How much water at 100°C can be changed to steam in 3mins by the heater? [Heat of vaporization= 2.26×10^6 J/kg, specific heat capacity of water = 4.2×10^3 J/kg.k].
Thank you
1
Expert's answer
2019-06-03T09:58:41-0400

From the conservation of energy:


"E=Pt=mL"

"m=\frac{Pt}{L}"

"m=\frac{(900)(3)(60)}{ 2.26\cdot 10^6 }=0.0717\ kg=71.7\ g"